OpenStudy (pokemon23):

anyone willing me to teach me how to simplify radicals been having problems with them

OpenStudy (lollylau):

you have any problems to do?

OpenStudy (akshay_budhkar):

you are not fluent with it?

OpenStudy (lollylau):

e.g. sqrt(8) =sqrt(4*2) =sqrt(4)sqrt(2) =2sqrt(2) e.g. 5rt(128) =5rt(32*4) =5rt(32)5rt(4) =2 5rt(4)

OpenStudy (pokemon23):

akshay can you help me

OpenStudy (pokemon23):

well can you assign me problems

OpenStudy (akshay_budhkar):

got 12-20 minutes, it is 00.45 here

OpenStudy (akshay_budhkar):

yea sure

OpenStudy (akshay_budhkar):

\[\sqrt 90 + \sqrt 48 - \sqrt 128\]

OpenStudy (pokemon23):

teach me tomorrow

OpenStudy (lollylau):

lol

OpenStudy (pokemon23):

it seems you need rest

OpenStudy (akshay_budhkar):

ok sure man :D meet me early

OpenStudy (mimi_x3):

still need help ? :/

OpenStudy (pokemon23):

I SAID TOMORROW @-@

OpenStudy (mimi_x3):

chilax dude..